Expected and required rates of return

Assume that the risk-free rate is 2.5% and the market risk premium is 6%.

What is the required return for the overall stock market? Round your answer to two decimal places.

What is the required rate of return on a stock with a beta of 0.6? Round your answer to two decimal places.
Answer #1

market risk premium =market rate-risk free rate

a.market rate=(2.5+6)=8.5%

b.Required return=risk free rate+beta*(market rate-risk free rate)

=2.5+0.6*(8.5-2.5)

=6.1%